OpenStudy (anonymous):

If f(x)= 4x+9 , find f(a+h)-f(a)/h

OpenStudy (anonymous):

first step is to find (carefully ) \[f(a+h)\] which will be \[4(a+h)+9=4a+4h+9\]

OpenStudy (anonymous):

from that, subtract \(f(a)=4a+h\) in other words, compute \[4a+4h+9-(4a+9)\]

OpenStudy (anonymous):

let me know what you get, there is only one more step after that

OpenStudy (anonymous):

do i have to distribute the - ?

OpenStudy (anonymous):

oh yes!

OpenStudy (anonymous):

for it to be 4a+4h+9-4a-9

OpenStudy (anonymous):

that is the first step yes, then do the actual subtraction, i.e. combine like terms

OpenStudy (anonymous):

so that means it is 4h ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and finally, divide that by \(h\) and what do you get?

OpenStudy (anonymous):

4:)

OpenStudy (anonymous):

yes, zactly this tells you that if \(f(x)\) is a line like \(f(x)=mx+b\) then \[\frac{f(a+h)-f(a)}{h}=m\] the slope

OpenStudy (anonymous):

aka "the change in y over the change in x" if \(f\) is not a line, then it is something else

OpenStudy (anonymous):

oh okay ! :D Thank you so much !:) @satellite73

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@satellite73 :)

OpenStudy (anonymous):

:) to you as well